The plates of a parallel-plate capacitor are 2.50 mm apart, and each carries a charge of


magnitude 80.0 nC. The plates are in vacuum. The electric field between the plates has a magnitude


of 4.00 X 106 V/m(a) What is the potential difference between the plates? (b) What is the area of


each plate? (c) What is the capacitance?



1
Expert's answer
2022-04-19T12:05:33-0400

a) The potential difference is:


"V = Ed"

where "E= 4.00 \times 10^6 V/m" and "d=2.50mm = 2.50\times 10^{-3}m". Thus, obtain:


"V = 4.00 \times 10^6 V/m\cdot 2.50\times 10^{-3}m = 1.00\times 10^4V"

c) The capacitance is given as follows:


"C = \dfrac{q}{V}"

where "q = 80.0\times 10^{-9}C". Obtain:


"C = \dfrac{80.0\times 10^{-9}C}{1.00\times 10^4V}=8.00\times 10^{-13} F"

b) On the other hand the capacitance of a parallel plate capacitor is:


"C = \dfrac{\varepsilon_0 A}{d}"

where "\varepsilon_0\approx 8.85\times 10^{-12} F/m" is the vacuum permittivity and "A" is the are of the plate. Expressing "A" and substituting "C", obtain:


"A = \dfrac{dC}{\varepsilon_0} = \dfrac{2.50\times 10^{-3}m\cdot 8.00\times 10^{-13} F}{8.85\times 10^{-12} F/m} \approx 2.26\times 10^{-4}m^2"